# An Information Theoretic Perspective on Conformal Prediction

### v1

@@ 1 Introduction
Machine Learning (ML) models have rapidly grown in popularity and reach, having now found use in many safety-critical domains like healthcare Ahsan et al. (2022) and autonomous driving Kuutti et al. (2020). In these areas, predictions must be accompanied by reliable measures of uncertainty to ensure safe decision-making, but unfortunately, most ML models are designed and trained to produce only point estimates, which capture only crude notions of uncertainty with no statistical guarantees. Conformal Prediction (CP) Vovk et al. (2005), in particular its split variant (SCP) Papadopoulos et al. (2002), has recently gained in popularity as a principled and scalable solution to equip any, potentially black-box, model with proper uncertainty estimates in the form of prediction sets; in loose terms, larger sets are associated with higher degrees of uncertainty.
In this work, we take a closer look at conformal prediction through the lens of Information Theory (IT), establishing a connection between conformal prediction and the underlying intrinsic uncertainty of the data-generating process, as captured by the conditional entropy $H(Y|X)$ of the target variable $Y$ given the inputfs $X$. We prove conformal prediction can be used to bound $H(Y|X)$ from above in three different ways: one derived from the data processing inequality, which we dub DPI bound, and two coming from a variation of Fano’s inequality Fano (1949), a model agnostic one, which we call (simple) Fano bound, and another informed by the predictive model itself, to which we refer as model-based Fano bound. To the best of our knowledge, these bounds represent the first bridge connecting information theory and conformal prediction, which we hope will bring new tools to both fields. We already present two such tools in this paper: (i) we show our upper bounds serve as principled training objectives to learn ML models that are more amenable to SCP, and (ii) we advance a systematic way to incorporate side information into the construction of prediction sets. We empirically validate that both these applications of our theoretical results lead to better predictive efficiency, i.e., narrower, more informative prediction sets.
@@ 2 Background
In this section we introduce the needed background on conformal prediction and list decoding (Elias, 1957; Wozencraft, 1958), a branch of information theory that, as we show, is closely related to conformal prediction and especially useful in deriving our main results. We start by introducing the notation used throughout the paper. As usual, we denote random variables in uppercase letters and their realization in lowercase, e.g., $X=x$. Similarly, we reserve calligraphic letters, e.g. ${\mathcal{X}}$, for sets. Throughout the paper, we use $P_{X},Q_{X},\dots$ to denote probability measures on the space ${\mathcal{X}}$. To simplify the notation, we simply use $P,Q,\dots$ when the underlying space is clear. For example, with a measure $Q_{XY}$ the probability of the event $\{(X,Y):Y\in{\mathcal{C}}(X)\}$ is denoted as $Q(Y\in{\mathcal{C}}(X))$.
@@ 2.1 Conformal Prediction
where the probability is over the randomness in the sample $\{(X_{i},Y_{i})\}_{i=1}^{n}\cup\{(X_{test},Y_{test})\}$. In this work, we focus on a variant called split conformal prediction (SCP) (Papadopoulos et al., 2002) that gained popularity in the ML community, since it can leverage any pre-trained model $f:{\mathcal{X}}\times{\mathcal{Y}}\rightarrow{\mathbb{R}}$ in the construction of prediction sets. In this setting, the aforesaid $n$ data points constitute a calibration data set ${\mathcal{D}}_{cal}$, which must be disjoint from the training data set used to fit the predictive model $f$. This separation between training and calibration data is what gives the name split to the method.
The first step in SCP is to define a score function $s_{f}:{\mathcal{X}}\rightarrow{\mathcal{Y}}$, which is itself a function of model $f$ and captures the magnitude of the prediction error at a given data point. We then evaluate the score function at every data point $(X_{i},Y_{i})\in{\mathcal{D}}_{cal}$ to obtain a collection of scores $\{S_{i}\}_{i=1}^{n}$, and at test time we construct prediction set ${\mathcal{C}}(X_{test})$ as
@@
$1-\alpha\leq\mathbb{P}(Y_{test}\in{\mathcal{C}}(X_{test}))\leq 1-\alpha_{n},\quad\text{ where}\quad\alpha_{n}=\alpha+\nicefrac{{1}}{{n+1}}.$
It is worth noting that valid coverage is not sufficient; the uninformative set predictor that always output ${\mathcal{C}}(X_{test})={\mathcal{Y}}$ trivially satisfies (1). We would also like our prediction sets to be as narrow as possible, and that is why CP methods are often compared in terms of their (empirical) inefficiency, i.e., the average prediction set size $\nicefrac{{1}}{{|D_{test}|}}\sum_{x\in{\mathcal{D}}_{test}}|{\mathcal{C}}(x)|$ for some test data set ${\mathcal{D}}_{test}$. This is, in fact, not the only type of inefficiency criterion (Vovk et al., 2016), but we use it as our main performance metric, since it is the most common.
@@ 2.2 Conformal Prediction as List Decoding
Define $\hat{y}=f(x)$. Upon observing $\hat{y}$, CP provides a list of candidates in the target set ${\mathcal{Y}}$. If we consider the mapping between $y$ and $\hat{y}$ as a noisy channel, the conformal prediction problem is equivalent to variable-size list decoding for the noisy channel between the true label $y$ and $\hat{y}$. In information theory and hypothesis testing, the decoder or the tester can provide a set of outcomes. If the correct solution is not part of the subset, an error is declared. The list decoding idea goes back to works of (Elias, 1957) and (Wozencraft, 1958) in communication theory followed by a plethora of works on the topic with various lower and upper bounds on the error. We review some of these results in Appendix B.2. In this paper, we leverage the information theoretic interpretation of CP mainly in two directions. First, we use the information theoretic inequalities for list decoding to upper bound the entropy of the data-generating process. This leads to new training mechanisms for conformal training (see Section 4). It additionally provides bounds on the inefficiency of a given model (see Appendix D and F.3). Second, we show we can incorporate side-information into CP to improve predictive efficiency (see Section 5).
@@ 3 Information Theory Applied to Conformal Prediction
In this section we develop our main results, linking information theory and conformal prediction. Concretely, we provide three different upper bounds on the conditional entropy $H(Y|X)$: one coming from the data processing inequality, and two derived in a similar way to Fano’s inequality.
@@ 3.1 Data Processing Inequality for Conformal Prediction
We can reuse some of the existing data processing inequalities (DPI) obtained for $f$-divergences to get bounds on conformal prediction. For a convex function $f$ with $f(1)=0$, the $f$-divergence between two probability measures $P$ and $Q$ is defined as (see Sason & Verdú (2016))
@@
In particular, with $f(x)=x\log x$ we recover the familiar notion of KL-divergence (Kullback & Leibler, 1951). The DPI for $f$-divergence states that for any two probability measures $P_{X}$ and $Q_{X}$ defined on a space ${\mathcal{X}}$, and any stochastic map $W_{Y|X}$, which maps $(P_{X},Q_{X})$ to $(P_{Y},Q_{Y})$, we have
$D_{f}(P_{X}||Q_{X})\geq D_{f}\left(P_{Y}||Q_{Y}\right),$
See Appendix B.1 for more details on $f$-divergence and data processing inequalities. The DPI would immediately yield the following proposition.
@@
where $h_{b}(\cdot)$ denotes the binary entropy function with $h_{b}(\alpha)=-\alpha\log(\alpha)-(1-\alpha)\log(1-\alpha).$
Note that the entropy term $H(Y|X)$ is computed using the measure $P$. We relegate the proof to Appendix C.1. In the above bound, the term $Q(Y\in{\mathcal{C}}(X))$ appears inside a log, so an empirical estimate $\hat{Q}(Y\in{\mathcal{C}}(X))$ would result in a lower bound and would be biased. We can provide an upper confidence bound on this estimate using the empirical Bernstein inequality Maurer & Pontil (2009), and use that instead. Based on the empirical Bernstein inequality, with probability $1-\delta$, we have
@@ 3.2 Model-Based Fano’s Inequality and Variations
Next, we present an inequality which is a variation of Fano’s inequality. Fano’s inequality is a classical result that, among other things, is used to prove the classical Shannon’s theorem on channel capacity. See Appendix B.4 for more details. We obtain the following inequality by modifying the classical proof of Fano’s inequality, which can be found in Cover & Thomas (2006).
@@
Note we have one term conditioned on the event of valid coverage, $Y\in{\mathcal{C}}(X)$, and another conditioned on $Y\notin{\mathcal{C}}(X).$ We provide the proof in Appendix C.2. A good choice for $Q$ is the predictive model itself, and that is why we refer to the bound above as Model-Based (MB) Fano bound. Another natural choice for $Q$ is the uniform distribution, in which case we get the following result.
@@
The proof follows directly from Proposition 3.2 by replacing $Q$ with the uniform distribution. We refer to the bound in (6) as (simple) Fano bound, since it is model agnostic and can be approximated directly with only empirical estimates of the prediction set size. This last bound explicitly relates the central notion of uncertainty in conformal prediction, the prediction set size, to an information theoretic concept of uncertainty in $H(Y|X)$. This information theoretic interpretation of conformal prediction as list decoding introduces various information theoretic tools, potentially useful for various applications. In Appendix D we derive some new inequalities for conformal prediction, in particular offering new lower bounds on the inefficiency of the conformal prediction. In the next section, we discuss how these inequalities can be used as conformal training schemes.
@@ 4 Conformal Training
Although we can apply split conformal prediction to any pretrained ML model as a post-processing step, the overall performance of any CP method (commonly measured by its inefficiency) is highly dependent on the underlying model itself. Therefore, previous works have proposed to take CP into account also during training and directly optimize for low predictive inefficiency (Bellotti, 2021; Stutz et al., 2022). We use the term conformal training to refer to such approaches in general (see Appendix E for an overview of the topic). In particular, we focus on ConfTr (Stutz et al., 2022) since it generalizes and outperforms (Bellotti, 2021).
In ConfTr (Stutz et al., 2022) each training batch ${\mathcal{B}}$ is split into calibration ${\mathcal{B}}_{cal}$ and test ${\mathcal{B}}_{test}$ halves to simulate the SCP process (see Section 2) for each gradient update of model $f$ and minimize the following size loss
@@
where ${\mathcal{C}}_{f}(x)$ is constructed using the statistics of ${\mathcal{B}}_{cal}$. Here we use the notation ${\mathcal{C}}_{f}(x)$ to make explicit the dependence of the prediction set on the model $f$. Still, SCP involves step functions and (Stutz et al., 2022) introduce a couple of relaxations to recover a differentiable objective: (i) the computation of quantiles is relaxed via differentiable sorting operators (Cuturi et al., 2019; Blondel et al., 2020; Petersen et al., 2022); (ii) the thresholding operation in the construction of prediction sets in (2) is replaced by smooth assignments of labels to prediction sets via the logistic sigmoid.
Our upper bounds on $H(Y|X)$, namely DPI (4), MB Fano (5) and simple Fano (6), presented in the previous section can be made differentiable in the same way, and thus can also serve as proper loss functions for conformal training. The motivation for doing so is twofold. First, the conditional entropy $H(Y|X)$ captures the underlying uncertainty of the task, or equivalently, the uncertainty under the true labelling distribution $P_{Y|X}$. Thus, by minimizing these upper bounds, we push the model $f$ closer to the true distribution, which is known to achieve minimal inefficiency (Vovk et al., 2016). Interestingly, the cross-entropy loss also bounds $H(Y|X)$ from above, and thus can be motivated as a conformal training objective from the same angle. In that regard, the DPI bound from Proposition 3.1 is particularly advantageous as it is provably tighter than the cross-entropy (see Appendix C.1).
Second, we can connect the simple Fano bound in (6) to the size loss (7) from (Stutz et al., 2022). In Appendix E.1, we show that via Jensen’s inequality and $\log(|Y|{-}|C(X)|)\leq\log|Y|$ the bound in (6) can be further upper bounded as
@@
$H(Y|X)\leq\lambda_{\alpha}+\left(1-\alpha_{n}\right)\log\mathbb{E}\left[|C(X)|\right],$
where $\lambda_{\alpha}$ only depends on $\alpha$ and $|{\mathcal{Y}}|$, and thus does not affect optimization. Therefore, we ground ConfTr as minimizing an upper bound to the true conditional entropy that is looser than the simple Fano bound, and likely also looser than the model-based Fano bound for an appropriate choice for $Q.$

### v2

@@ 1 Introduction
Machine learning (ML) models have rapidly grown in popularity and reach, having now found use in many safety-critical domains like healthcare [1] and autonomous driving [26]. In these areas, predictions must be accompanied by reliable measures of uncertainty to ensure safe decision-making. However, most ML models are designed and trained to produce only point estimates, which capture only crude notions of uncertainty with no statistical guarantees. Conformal prediction (CP) [62], in particular its split variant (SCP) [45], has recently gained in popularity as a principled and scalable solution to equip any, potentially black-box, model with proper uncertainty estimates in the form of prediction sets; in loose terms, larger sets are associated with higher degrees of uncertainty.
In this work, we take a closer look at conformal prediction through the lens of information theory (IT), establishing a connection between conformal prediction and the underlying intrinsic uncertainty of the data-generating process, as captured by the conditional entropy $H(Y|X)$ of the target variable $Y$ given the inputs $X$. We prove conformal prediction can be used to bound $H(Y|X)$ from above in three different ways: one derived from the data processing inequality, which we dub DPI bound, and two coming from a variation of Fano’s inequality [18], a model agnostic one, which we call (simple) Fano bound, and another informed by the predictive model itself, to which we refer as model-based Fano bound. To the best of our knowledge, these bounds represent the first bridge connecting information theory and conformal prediction, which we hope will bring new tools to both fields. We already present two such tools in this paper: (i) we show our upper bounds serve as principled training objectives to learn machine learning models that are more amenable to SCP, and (ii) we advance a systematic way to incorporate side information into the construction of prediction sets. We empirically validate that both these applications of our theoretical results lead to better predictive efficiency, i.e. narrower and, consequently, more informative prediction sets.
@@ 2 Background
In this section, we present the needed background on conformal prediction and list decoding [17, 66], an area of information theory that, as we show, is closely related to CP and especially useful in deriving our main results. We start with the necessary notation. As usual, we denote random variables in uppercase letters and their realization in lowercase, e.g., $X=x$. We reserve calligraphic letters, e.g. ${\mathcal{X}}$, for sets and use $P_{X},Q_{X},\dots$ to denote probability measures on the space ${\mathcal{X}}$. To simplify the notation, we use $P,Q,\dots$ when the underlying space is clear. For example, given a probability measure $Q_{XY}$, the probability of the event $\{(X,Y):Y\in{\mathcal{C}}(X)\}$ is denoted as $Q(Y\in{\mathcal{C}}(X))$.
@@ 2.1 Conformal Prediction
where the probability is over the randomness in the sample $\{(X_{i},Y_{i})\}_{i=1}^{n}\cup\{(X_{test},Y_{test})\}$. To make this more tangible, the reader can picture ${\mathcal{C}}(X)$ as a subset of the possible labels in a classification problem, or as a confidence interval around the point estimate of a regressor in a regression setting.
In this work, we focus on a variant called split conformal prediction (SCP) [45] that gained popularity in the ML community, since it can leverage any pre-trained model $f:{\mathcal{X}}\rightarrow{\mathcal{Y}}$ in the construction of prediction sets. In this setting, the aforesaid $n$ data points constitute a calibration data set ${\mathcal{D}}_{cal}$, which must be disjoint from the training data set used to fit the predictive model $f$. This separation between training and calibration data is what gives the name split to the method.
The first step in SCP is to define a nonconformity score function $s_{f}:{\mathcal{X}}\times{\mathcal{Y}}\rightarrow{\mathbb{R}}$, which is itself a function of model $f$ and captures the magnitude of the prediction error at a given data point; the higher the score $s_{f}(x,y)$, the higher the disagreement between input $x$ and prediction $y$. At calibration time, evaluate the score function at every $(X_{i},Y_{i})\in{\mathcal{D}}_{cal}$ to get a collection of scores $\{S_{i}=s_{f}(X_{i},Y_{i})\}_{i=1}^{n}$, and at test time, we construct prediction set ${\mathcal{C}}(X_{test})$ as
@@
$1-\alpha\leq\mathbb{P}(Y_{test}\in{\mathcal{C}}(X_{test}))\leq 1-\alpha_{n},\quad\text{ where}\quad\alpha_{n}=\alpha-\nicefrac{{1}}{{n+1}}.$
See Appendix B for the proof and a more thorough introduction to CP. It is worth noting that valid coverage is not sufficient; the uninformative set predictor that always outputs ${\mathcal{C}}(X_{test})={\mathcal{Y}}$ trivially satisfies (1). We would also like our prediction sets to be as narrow as possible, and that is why CP methods are often compared in terms of their (empirical) inefficiency, i.e., the average prediction set size $\nicefrac{{1}}{{|D_{test}|}}\sum_{x\in{\mathcal{D}}_{test}}|{\mathcal{C}}(x)|$ for some test data set ${\mathcal{D}}_{test}$. This is, in fact, not the only type of inefficiency criterion, but we use it as our main performance metric since it is the most common [63].
@@ 2.2 Conformal Prediction as List Decoding
In a nutshell, we can see conformal prediction as defining a map from an input $x$ to a set of candidates in the target set ${\mathcal{Y}}$. It turns out that the conformal prediction framework is equivalent to (variable-size) list decoding, an error-recovery model going back to the works of Elias [17] and Wozencraft [66] in communication theory—we review some of these results in Appendix C.2. In particular, consider the mapping from the true label $y$ to the input $x$ as a noisy communication channel $p(x|y)$. The goal of an error-correcting code is then to decode the input $x$ and recover the one true label $y$. List decoding generalizes this idea, allowing the decoder to return a set of outcomes (a list) instead of a point-wise prediction. If the correct solution is not part of the set output by the decoder, an error is declared. Although conformal prediction and list decoding were developed for different purposes, namely uncertainty quantification and error correction, it is easy to see that, if we allow for variable-size lists, the list decoding problem for the channel $p(x|y)$ as described above is equivalent to the conformal prediction problem.
To our knowledge, this link between conformal prediction and information theory (and list decoding in particular) has gone unnoticed in the literature, and in this paper we leverage it in two directions. First, we apply information-theoretic inequalities for list decoding to upper bound the conditional entropy $H(Y|X)$ of the data-generating process. This leads to new objectives for conformal training (see Section 4) and new bounds on the inefficiency of a given model (see Appendices E and G.3). Second, the information-theoretic interpretation of CP gives us an effective and theoretically grounded way of incorporating side-information into CP to improve predictive efficiency (see Section 5).
@@ 3 Information Theory Applied to Conformal Prediction
In this section, we develop our main results, which link information theory and conformal prediction. Concretely, we provide three novel upper bounds on the conditional entropy $H(Y|X)$: one coming from the data processing inequality and two derived in a similar way to Fano’s inequality. We defer the proofs to Appendix D for the sake conciseness, but in broad strokes, our results come from relating the bounds on the error probability provided by these information-theoretic inequalities (typically in the context of error-correcting codes) to the guarantees provided by CP in Theorem 2.1.
@@ 3.1 Data Processing Inequality for Conformal Prediction
We start by using the classical data processing inequality (DPI) in the context of conformal prediction. Specifically, we focus on the DPI for $f$-divergences, which discuss thoroughly in Appendix C.1. In brief, for a convex function $f$ with $f(1)=0$, the $f$-divergence between two probability measures $P$ and $Q$ is defined as (see [56])
@@
In particular, with $f(x)=x\log x$ we recover the familiar notion of KL-divergence [25]. The DPI for $f$-divergences states that for any two probability measures $P_{X}$ and $Q_{X}$ defined on a space ${\mathcal{X}}$, and any stochastic map $W_{Y|X}$, which maps $(P_{X},Q_{X})$ to $(P_{Y},Q_{Y})$, we have
$D_{f}(P_{X}||Q_{X})\geq D_{f}\left(P_{Y}||Q_{Y}\right).$
We can apply the DPI for $f$-divergence above in the context of conformal prediction by considering the probability of the event of valid coverage $\{Y\in{\mathcal{C}}(X)\}$ under two different probability measures $P$ and $Q$. Taking $P$ as the data-generating distribution $P:=P_{X}P_{Y|X}$ and constructing $Q:=P_{X}Q_{Y|X}$ for an arbitrary $Q_{Y|X}$ (e.g., a machine learning model), we get the following proposition.
@@
with $\alpha_{n}=\alpha-\nicefrac{{1}}{{n+1}}$ and $h_{b}(\cdot)$ the binary entropy function $h_{b}(\alpha)=-\alpha\log(\alpha)-(1-\alpha)\log(1-\alpha).$
Note that the entropy term $H(Y|X)$ is computed using the measure $P$. We relegate the proof to Appendix D.1. In the bound in (3), the term $Q(Y\in{\mathcal{C}}(X))$ appears inside a log, so an empirical estimate $\hat{Q}(Y\in{\mathcal{C}}(X))$ would result in a lower bound and would be biased. We can provide an upper confidence bound on this estimate using the empirical Bernstein inequality [39], and use that instead. Based on the empirical Bernstein inequality, with probability $1-\delta$, we have
@@ 3.2 Model-Based Fano’s Inequality and Variations
Next, we present an inequality which is a variation of Fano’s inequality [18], a classical result that, among other things, is used to prove Shannon’s classical theorem on channel capacity. See Appendix C.4 for more details. In our context, we can use Fano’s inequality to relate the conditional entropy $H(Y|X)$ to the probability of error, i.e., $\mathbb{P}(Y\notin{\mathcal{C}}(X))$. From that insight, we obtain Proposition 3.2 by modifying the classical proof of Fano’s inequality, which can be found in [13], and applying the conformal guarantees from Theorem 2.1 to bound the probability of error.
@@
Note that we have one term conditioned on the event of valid coverage, $\{Y\in{\mathcal{C}}(X)\}$, and another conditioned on $\{Y\notin{\mathcal{C}}(X)\}.$ We provide the proof in Appendix D.2. A good choice for $Q$ is the predictive model itself, and that is why we refer to the bound above as Model-Based (MB) Fano bound. Another natural choice for $Q$ is the uniform distribution, which gives us the following result.
@@
The proof follows directly from Proposition 3.2 by replacing $Q$ with the uniform distribution. We refer to the bound in (6) as (simple) Fano bound, since it is model agnostic and can be approximated directly with only empirical estimates of the prediction set size. This last bound explicitly relates the central notion of uncertainty in conformal prediction, the prediction set size, to an information-theoretic concept of uncertainty in $H(Y|X)$. This reinterpretation of conformal prediction as a form of list decoding introduces various information-theoretic tools, potentially useful for various applications. In Appendix E we derive some new inequalities for conformal prediction, in particular offering new lower bounds on the inefficiency of the conformal prediction. In the next section, we discuss how these inequalities can be used as conformal training schemes.
@@ 4 Conformal Training
Although split conformal prediction is applicable to any pretrained ML model as a post-processing step, the overall performance of any CP method (commonly measured by its inefficiency) is highly dependent on the underlying model itself. Therefore, previous works have proposed to take CP into account already during model training and directly optimize for low predictive inefficiency [8, 59]. We use the term conformal training to refer to such approaches in general (see Appendix F for an overview of the topic). In particular, we focus on ConfTr [59] since it generalizes and outperforms [8]. In ConfTr [59] each training batch ${\mathcal{B}}$ is split into calibration ${\mathcal{B}}_{cal}$ and test ${\mathcal{B}}_{test}$ halves to simulate the SCP process (see Section 2) for each gradient update of model $f$ and minimize the following size loss
@@
where ${\mathcal{C}}_{f}(x)$ is constructed using the statistics of ${\mathcal{B}}_{cal}$. Here we use the notation ${\mathcal{C}}_{f}(x)$ to make explicit the dependence of the prediction set on the model $f$. Still, SCP involves step functions and Stutz et al. [59] introduce a couple of relaxations to recover a differentiable objective: (i) the computation of quantiles is relaxed via differentiable sorting operators [16, 9, 48]; (ii) the thresholding operation in the construction of prediction sets in (2) is replaced by smooth assignments of labels to prediction sets via the logistic sigmoid.
Our upper bounds on $H(Y|X)$, namely DPI, MB Fano and simple Fano, presented in the previous section can be made differentiable in the same way and thus can also serve as proper loss functions for conformal training. The motivation for doing so is twofold. First, the conditional entropy $H(Y|X)$ captures the underlying uncertainty of the task, or equivalently, the uncertainty under the true labelling distribution $P_{Y|X}$. Thus, by minimizing these upper bounds, we can hope to push the model $f$ closer to the true distribution, which is known to achieve minimal inefficiency [63]. Interestingly, the cross-entropy loss also bounds $H(Y|X)$ from above and thus can be motivated as a conformal training objective from the same angle. In that regard, the DPI bound from Proposition 3.1 is particularly advantageous as it is provably tighter than the cross-entropy (see Appendix D.1).
Second, we can connect the simple Fano bound from Corollary 3.1 to the size loss (7) from [59]. In Appendix F.1, we show that via Jensen’s inequality and $\log(|Y|{-}|C(X)|)\leq\log|Y|$ the bound in Corollary 3.1 can be further upper bounded as
@@
$H(Y|X)\leq\lambda_{\alpha}+\left(1-\alpha_{n}\right)\log\mathbb{E}\left[|C(X)|\right],$ (8)
Since $\lambda_{\alpha}$ and $(1-\alpha_{n})$ are constants, they do not affect optimization, and minimizing the right hand side in (4) is equivalent to minimizing the size loss in (7). Therefore, we ground ConfTr as minimizing an upper bound to the true conditional entropy that is looser than the simple Fano bound and likely also looser than the model-based Fano bound for an appropriate choice for $Q.$